Find |u| if u=<-6,6>
Answer by stanbon(75887)

|u| if u=<-6,6>
------------
|u| = sqrt(6^2 + 6^2) = sqrt(2*36) = 6*sqrt(2)
-----------------
Cheers,
Stan H.
---------------